# Multi-Sensor Diffusion-Driven Optical Image Translation for Large-Scale Applications

### v3

@@ Abstract
Comparing images captured by disparate sensors is a common challenge in remote sensing. This requires image translation—converting imagery from one sensor domain to another while preserving the original content. Denoising Diffusion Implicit Models (DDIM) are potential state-of-the-art solutions for such domain translation due to their proven superiority in multiple image-to-image translation tasks in classic computer vision. However, these models struggle with large-scale multi-patch imagery, often focusing solely on small patches and resulting in inconsistencies across the full image. To overcome these limitations, we propose a novel method that leverages DDIM for effective optical image translation over large areas. Our approach super-resolves large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. Our contributions lie in new forward and reverse diffusion processes, which are crucial for addressing the challenges of large-scale image translation. Extensive experiments with a dataset of paired Sentinel-II and Planet Dove images demonstrate that our approach provides precise domain adaptation and artifact reduction. The proposed method preserves image content while improving radiometric (color) accuracy and feature representation. The outcome is a high-resolution large-scale image with consistent patches, vital for applications such as heterogeneous change detection (HCD). We present a unique training and testing algorithm rooted in DDIMs, a thorough image quality assessment, and a comparative study against the standard classifier-free guided DDIM framework and five other leading methods. The usefulness of our approach is further demonstrated in HCD tasks performed in the urban settings of Beirut, Lebanon, and Austin, USA.
@@ I Introduction
Despite advancements in image-to-image (I2I) translation for remote sensing, most literature focuses on generating individual synthetic patches for augmenting training data sets used in downstream tasks [I2ITasar2020, I2IIsmael2023, I2ISokolov2023], or on SAR to Optical translation [CDLi2021, I2IZhang2023, I2ISeo2023]. While highly valuable, these methods do not address the challenge of translating large-scale images across disparate optical sensors with different spatial resolutions, particularly ensuring feature consistency between input and output images. This gap is significant as it hinders the fidelity of the translated images, where preserving salient features without introducing non-existent elements is crucial. Additionally, these methods are usually tested on benchmark datasets or small datasets, which may not adequately demonstrate their effectiveness in real-world scenarios.
@@ II-1 Denoising Diffusion Models
Ho et al. [Ho2020] proposed the use of a denoising framework for image generation, the so-called Denoising Diffusion Probabilistic Model (DDPM). It showed incredible image generation quality and overcame most state-of-the-art methods available at the time. However, the inference process was cumbersome since hundreds of model evaluations per image were necessary to reach optimal performance. The work Denoising Diffusion Implicit Models (DDIM) [Song2020] approached this issue by presenting a new interpretation of the reverse diffusion process. Differently from the work in [Ho2020], they treated reverse diffusion as a non-Markovian process by skipping multiple reverse diffusion steps. Furthermore, they experimented with eliminating the randomness of intermediate diffusion steps—which came from the addition of a random noise matrix to each reverse diffusion step result—and noticed that this enabled an even greater reduction of model evaluations during inference at the cost of sample variety. In practice, these combined modifications reduced the necessary number of model evaluations during inference by up to two orders of magnitude while maintaining high-generation quality.
@@
As originally defined in [Ho2020], denoising diffusion models are unconditional, i.e., they are trained to generate images with distributions similar to those it has seen during training without any prior knowledge about how a particular image should resemble or what it should contain. This limited the applicability of such models in multiple condition-guided tasks, such as image inpainting, super-resolution, colorization, and domain adaptation. To make them possible, two works [Nichol2021, Saharia2021] experimented with adding a conditioning matrix as an additional model input. This partially worked since, as observed by Dhariwal et al. [Dhariwal2021], diffusion models have a tendency to ignore the condition and focus only on the noisy input to be denoised. The work in [Dhariwal2021] also proposed a possible strategy to alleviate this issue: the use of a separate classifier through which conditioning information is directly inserted into the reverse diffusion procedure. This proved to be useful, as the images generated with this procedure were strongly linked with the conditions added to the model. Still, the need to have a second classifier model is cumbersome and sometimes impractical. Due to that, Ho et al. [ho2022classifierfree] proposed Classifier-free Guidance, which, as the name implies, removes the need for a separate classifier network that guides the model with conditioning information. They accomplish that by making a small modification in the DDPM training and inference procedures that, in summary, treats the diffusion model as if it consisted of a combination of an unconditional and a conditional model. Hence, during inference, it became possible to determine how much the conditional and unconditional virtual parts of the model influenced the generation and, therefore, how strongly the reverse diffusion should rely on the conditioning information.
@@ II-2 Image-to-Image Translation
Other works focused on remote sensing-related tasks. Ismael et al. [I2IIsmael2023] proposed a GAN-based I2I translation method with the goal of performing remote sensing semantic segmentation of optical images from unknown domains using a segmentation network trained with optical images from a single domain. The methods in [I2ITasar2020] and [I2ISokolov2023] proposed style-transferring domain adapting frameworks that generate extra training data matching the appearance of a particular same-resolution domain to alleviate train/test data drift. The method in [CDLi2021] uses a cycle-consistent GAN to translate optical images into SAR images to perform change detection. It showed promising change detection results. However, training and testing data are limited to just a few acquisitions. The method in [I2IZhang2023] is a thermodynamics-inspired translation network that obtained interesting results in SAR-to-Optical translation tasks with its unconventional design. The work in [I2ISeo2023] proposes a DDM-based Brownian Bridge method for SAR-to-Optical I2I translation. There, the authors suggest the use of a modified forward diffusion procedure that, instead of resulting in a pure-noise image, the target SAR image is iteratively transformed into the source optical image. They reason that this could help the model more easily understand the relationships across domains. Xiao et al. propose a diffusion-based super-resolution method that seeks to further explore prior information from the low-resolution input. When tested on a small-scale data set, it performed well.
@@ III Proposed Method
The method hereby proposed exclusively employs Denoising Diffusion Implicit Models (DDIM) [Song2020], which are advantageous due to their reduced number of model forward passes during inference and their deterministic behavior in intermediate steps. However, our approach differs from the standard classifier-free guided DDIM framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022]. The standard framework struggles with the high variability in patch coloration within large training sets, leading to inaccurate colorization in generated patches.
@@ III-A Training
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the similar features. Thus, patches aligned in time are required for training to work. Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$. The matrix $\bm{I}_{\text{in}}^{\text{LR,local}}$ is defined as the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference. $\bm{I}_{\text{in}}^{\text{LR, global}}$ represents the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast patches that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. This is done to give more context to the model of what exactly lies around the input patch. Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix (like in [ho2022classifierfree]). To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until convergence.
@@ V-A Image Quality Metrics and Ablation Experiments
Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting (Proposed) 0.1884 45.64 15.72
@@
SR3 [Saharia2021SR3] 0.6235 152.6 9.906
1 Identical neural network architecture as the proposed method.
@@
We compare our proposed I2I algorithm with other possible I2I solutions. First, the Sentinel-II test set is directly compared to Planet Dove images without any I2I translation method. The obtained metrics are shown in the first row of Table I. Second, we train two deep regression models [Lathuiliere2020] whose loss objective is to directly match the pixel values of Planet patches when Sentinel-II corresponding patches are input. They share the same neural network backbone as the proposed DDM-based method. In other words, these models are identical to the one used in the proposed method, with the exception that the output linear layer directly predicts the values of pixels of the corresponding Planet patch for a given input. The first deep regression model, whose metrics are displayed in the second row of Table I, is a conventional regression model, i.e., it does not apply any pre or post-processing to its inputs and outputs apart from simple $[-1,1]$ normalization. The second regression model, whose results are displayed in the third row, is identical to the previous model, except that the whitening and coloring algorithms, defined in Algorithm 2 and 4, are applied to its inputs and outputs, respectively. In the table, we refer to this as Regression + $\mathbf{W},\mathbf{C}$. The goal of testing this configuration is to see what would be the influence of these processing methods in a conventional regression model and whether or not it has a significant impact on the metrics. Following, we test the standard classifier-free guided DDIM [ho2022classifierfree], i.e., without the proposed whitening, coloring, and PSNR voting processes, using the exact same model architecture as the proposed method, to perform an ablation of these techniques. It is referred in Table I as Conditional DDIM. Then, we train an identical model, now adding the whitening and coloring procedures. To measure the impact of the PSNR voting in the inference results, the image quality metrics for this model are separated into two rows, the first being without PSNR voting—denoted as Conditional DDIM + $\mathbf{W},\mathbf{C}$—and the second with it—Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. All DDIM results were obtained using $N_{\text{final}}=64$. The proposed method, which includes PSNR voting, used $N_{\text{pre}}=8$ and $n_{\text{noisy}}=8$.
@@
We also compare the performance of our method with five other state-of-the-art approaches for super resolution and image-to-image translation tasks. The first method is the well-established Pix2Pix [Pix2Pix2017], which has been proposed for multiple generative tasks, including image-to-image translation. It is an adversarial framework that uses both regression and discriminative losses to guide training of a generator UNet-like network. Following, we compare our method against the SwinIR [liang2021swinir] super resolution model, which leverages the Swin Transformer architecture for enhanced image processing capabilities. The ShuffleMixer [sun2022shufflemixer] architecture, known for its efficient mixing of features through channel shuffling and spatial mixing, is another model we compare against. The SRDenseNet [zhang2018residual] is also included in our comparison—a model that utilizes dense connections to improve the flow of information and gradients throughout the network, which is particularly beneficial for the task of super resolution. Finally, we present results for the SR3 super resolution diffusion-based framework [Saharia2021SR3], using DDIM sampling with 64 steps. The UNet structure of SR3 is analogous to the one used in the article that introduced the DDPM procedure [Ho2020]. Thus, comparing it to our method provides an insight on the advantages of the neural network structure chosen for our experiments.

### v4

@@ Abstract
Comparing images captured by disparate sensors is a common challenge in remote sensing. This requires image translation—converting imagery from one sensor domain to another while preserving the original content. Denoising Diffusion Implicit Models (DDIM) are potential state-of-the-art solutions for such domain translation due to their proven superiority in multiple image-to-image translation tasks in computer vision. However, these models struggle with reproducing radiometric features of large-scale multi-patch imagery, resulting in inconsistencies across the full image. This renders downstream tasks like Heterogeneous Change Detection impractical. To overcome these limitations, we propose a method that leverages denoising diffusion for effective multi-sensor optical image translation over large areas. Our approach super-resolves large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. Our contributions lie in new forward and reverse diffusion processes that address the challenges of large-scale image translation. Extensive experiments using paired Sentinel-II (10m) and Planet Dove (3m) images demonstrate that our approach provides precise domain adaptation, preserving image content while improving radiometric accuracy and feature representation. A thorough image quality assessment and comparisons with the standard DDIM framework and five other leading methods are presented. We reach a mean Learned Perceptual Image Patch Similarity (mLPIPS) of 0.1884 and a Fréchet Inception Distance (FID) of 45.64, expressively outperforming all compared methods, including DDIM, ShuffleMixer, and SwinIR. The usefulness of our approach is further demonstrated in two Heterogeneous Change Detection tasks.
@@ I Introduction
Despite advancements in multi-sensor image-to-image (I2I) translation for remote sensing, most literature focuses on generating individual synthetic patches for augmenting training data sets used in downstream tasks [I2ITasar2020, I2IIsmael2023, I2ISokolov2023], on SAR to Optical translation [CDLi2021, I2IZhang2023, I2ISeo2023], or in super-resolution without tackling multisensor discrepancies beyond resolution differences [yi_tip2024, Min2024, Haut2018]. While highly valuable, these methods do not address the challenge of translating large-scale images across disparate optical sensors with different spatial resolutions, particularly ensuring feature consistency between input and output images and preserving the radiometric characteristics of each sensor. This gap is significant as it hinders the fidelity of the translated images, where preserving salient features without introducing non-existent elements is crucial. Additionally, these methods are usually tested on benchmark datasets or small datasets, which may not adequately demonstrate their effectiveness in real-world scenarios.
@@ II-1 Denoising Diffusion Models
Ho et al. [Ho2020] proposed the use of a denoising framework for image generation, the so-called Denoising Diffusion Probabilistic Model (DDPM). It showed incredible image generation quality and overcame most state-of-the-art methods available at the time. However, the inference process was cumbersome since hundreds of model evaluations per image were necessary to reach optimal performance. The work Denoising Diffusion Implicit Models (DDIM) [Song2020] approached this issue by presenting a new interpretation of the reverse diffusion process. Differently from the work in [Ho2020], they treated reverse diffusion as a non-Markovian process by skipping multiple reverse diffusion steps. Furthermore, they experimented with eliminating the randomness of intermediate diffusion steps—which came from the addition of a random noise matrix to each reverse diffusion step result—and noticed that this enabled an even greater reduction of model evaluations during inference at the cost of sample variety. In practice, these combined modifications reduced the necessary number of model evaluations during inference by up to two orders of magnitude while maintaining high generation quality.
@@
As originally defined in [Ho2020], denoising diffusion models are unconditional, i.e., they are trained to generate images with distributions similar to those it has seen during training without any prior knowledge about how a particular image should resemble or what it should contain. This limited the applicability of such models in multiple condition-guided tasks, such as image inpainting, super-resolution, colorization, and domain adaptation. To make them possible, two works [Nichol2021, Saharia2021] experimented with adding a conditioning matrix as an additional model input. This partially worked since, as observed by Dhariwal et al. [Dhariwal2021], diffusion models have a tendency to ignore the condition and focus only on the noisy input to be denoised. Furthermore, the work in [Dhariwal2021] proposed a possible strategy to alleviate this issue: the use of a separate classifier through which conditioning information is directly inserted into the reverse diffusion procedure. This proved to be useful, as the images generated with this procedure were strongly linked with the conditions added to the model. Still, the need to have a second classifier model is cumbersome and sometimes impractical. Due to that, Ho et al. [ho2022classifierfree] proposed Classifier-free Guidance, which, as the name implies, removes the need for a separate classifier network that guides the model with conditioning information. They accomplish that by making a small modification in the DDPM training and inference procedures that, in summary, treats the denoising model as if it consisted of a combination of an unconditional and a conditional model. Hence, during inference, it became possible to determine how much the conditional and unconditional virtual parts of the model influenced the generation and, therefore, how strongly the reverse diffusion should rely on the conditioning information.
@@ II-2 Image-to-Image Translation
Other works focused on remote sensing-related tasks. Ismael et al. [I2IIsmael2023] proposed a GAN-based I2I translation method with the goal of performing remote sensing semantic segmentation of optical images from unknown domains using a segmentation network trained with optical images from a single domain. The methods in [I2ITasar2020] and [I2ISokolov2023] proposed style-transferring domain adapting frameworks that generate extra training data matching the appearance of a particular same-resolution domain to alleviate train/test data drift. The method in [CDLi2021] uses a cycle-consistent GAN to translate optical images into SAR images to perform change detection. It showed promising change detection results. However, training and testing data are limited to just a few acquisitions. The method in [I2IZhang2023] is a thermodynamics-inspired translation network that obtained interesting results in SAR-to-Optical translation tasks with its unconventional design. The work in [I2ISeo2023] proposes a DDM-based Brownian Bridge method for SAR-to-Optical I2I translation. There, the authors suggest the use of a modified forward diffusion procedure that, instead of resulting in a pure-noise image, the target SAR image is iteratively transformed into the source optical image. They reason that this could help the model more easily understand the relationships across domains. Xiao et al. [Xiao2023] propose a diffusion-based super-resolution method that seeks to further explore prior information from the low-resolution input. When tested on a small-scale data set, it performed well.
In the context of remote sensing super-resolution, [yi_tip2024] focuses on improving image super-resolution techniques in large-area earth observation scenes by addressing issues with redundant token representation and single-scale modeling. They introduce the Top-k Token Selective Transformer (TTST), which uses a Residual Token Selective Group (RTSG) and a Multi-scale Feed-forward Layer (MFL) for better feature aggregation and accurate reconstruction. The TTST outperforms existing methods without attempting multi-sensor translation. Min et al. [Min2024] introduce a domain matching (DM) module that enhances reference-based image super-resolution (RefSR) models by addressing domain gaps between reference and low-resolution images in remote sensing. This DM-based RefSR approach outperformed the compared methods. However, its reliance on reference image features can potentially introduce unwanted artifacts not present in the input image, and it requires co-registered reference images with similar geometry of acquisition, limiting its practical applicability. Haut et al. [Haut2018] present an unsupervised convolutional generator model for super-resolving low-resolution remote sensing data. Data is symmetrically projected to the target resolution while ensuring reconstruction constraints over the LR input image. This study did not attempt multi-sensor translation, but rather downsampled high-resolution images to create low-resolution counterparts.
@@ III Proposed Method
The method hereby proposed is built upon Denoising Diffusion Implicit Models (DDIM) [Song2020], which are advantageous due to their reduced number of model forward passes during inference and their deterministic behavior in intermediate steps. However, our approach differs from the standard classifier-free guided DDIM framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022]. The standard framework struggles with the high variability in patch coloration within large training sets, leading to inaccurate colorization in generated patches. This color shifting behavior in diffusion models has been originally described by Deck and Bischoff [deck2023easing]. They observed that diffusion models can exhibit color shifts that are worsened with larger input patch sizes.
@@ III-A Training
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the similar features. Thus, patches aligned in time are required for training to work. Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$. The matrix $\bm{I}_{\text{in}}^{\text{LR,local}}$ is defined as the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference. $\bm{I}_{\text{in}}^{\text{LR, global}}$ represents the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast patches that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. That is, $\bm{I}_{\text{in}}^{\text{LR,local}}$ is the southwestern patch in a 2x2 patch grid. This grid is downsampled to the pixel dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$ to form $\bm{I}_{\text{in}}^{\text{LR, global}}$. This is done to give more context to the model of what exactly lies in the vicinity the input patch. Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix (like in [ho2022classifierfree]). To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until convergence.
@@ V-A Image Quality Metrics and Ablation Experiments
Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting (Proposed) 1 0.1884 45.64 15.72
@@
Diffusion Using the SR3 Backbone [Saharia2021SR3]
Conditional DDIM 0.6235 152.6 9.906
Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting 0.3730 82.15 15.87
1Using the SDM [Wang2022] backbone.
@@
We compare our proposed I2I algorithm, which has been trained using the SDM [Wang2022] neural network as backbone, with other possible I2I solutions. First, the Sentinel-II test set is directly compared to Planet Dove images without any I2I translation method. The obtained metrics are shown in the first row of Table I. Second, we train two deep regression models [Lathuiliere2020] whose loss objective is to directly match the pixel values of Planet patches when Sentinel-II corresponding patches are input. They share the same neural network backbone as the proposed DDM-based method. These models are identical to the one used in our experiments, except that the output linear layer directly predicts the pixel values of the corresponding Planet patch for a given input. The first deep regression model, whose metrics are displayed in the second row of Table I, is a conventional regression model without any pre- or post-processing apart from simple $[-1,1]$ normalization. The second regression model, shown in the third row, is identical but includes the whitening and coloring algorithms defined in Algorithms 2 and 4. In the table, we refer to this as Regression + $\mathbf{W},\mathbf{C}$. Testing this configuration assesses the influence of these processing methods in a conventional regression model and their impact on the metrics. Next, we test the standard classifier-free guided DDIM [ho2022classifierfree], i.e., without the proposed whitening, coloring, and PSNR voting processes, using the exact same model architecture as the proposed method, to perform an ablation of these techniques. It is referred to in Table I as Conditional DDIM. We then train an identical model, now adding the whitening and coloring procedures. To measure the impact of the PSNR voting in the inference results, the image quality metrics for this model are separated into two rows, the first being without PSNR voting—denoted as Conditional DDIM + $\mathbf{W},\mathbf{C}$—and the second with it—Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. All DDIM results were obtained using $N_{\text{final}}=64$. The proposed method, which includes PSNR voting, used $N_{\text{pre}}=8$ and $n_{\text{noisy}}=8$.
@@
We also compare the performance of our method with five other state-of-the-art approaches for super resolution and image-to-image translation tasks. The first method is the well-established Pix2Pix [Pix2Pix2017], which has been proposed for multiple generative tasks, including image-to-image translation. It is an adversarial framework that uses both regression and discriminative losses to guide training of a generator UNet-like network. Following, we compare our method against the SwinIR [liang2021swinir] super resolution model, which leverages the Swin Transformer architecture for enhanced image processing capabilities. The ShuffleMixer [sun2022shufflemixer] architecture, known for its efficient mixing of features through channel shuffling and spatial mixing, is another model we compare against. The SRDenseNet [zhang2018residual] is also included in our comparison—a model that utilizes dense connections to improve the flow of information and gradients throughout the network, which is particularly beneficial for the task of super resolution. Finally, we present results for the SR3 super resolution diffusion-based framework [Saharia2021SR3]. The UNet structure of SR3 is analogous to the one used in the article that introduced the DDPM procedure [Ho2020]. Thus, comparing it to our method provides an insight on the advantages of the neural network structure chosen for our experiments.